Calculate the mass of Earth. Given, the force that Earth exerts on a
mass (m) of 1 kg at its surface is 9.8 N. The radius of Earth ( R ) is
6. 4 × 10⁶ m.
1
Expert's answer
2020-11-30T15:22:24-0500
"g=\frac{GM}{R^2}\\\frac{(6.67\cdot10^{-11})}{(6.4\cdot10^{6})^2}M=9.8\\M=6.0\cdot10^{24}kg"